OpenStudy (anonymous):

Determine whether the point (2, 0) is a solution to the system of equations. Explain your reasoning in complete sentences.

OpenStudy (sohailiftikhar):

just put these point in the place of variable and if they satisfy the equeation then they are solution sets ......

OpenStudy (anonymous):

how do I know if they satisfy them? I plugged it in and i got g(x) = 8 and f(x) = 2

OpenStudy (anonymous):

That is a picture of the graph they gave us

OpenStudy (sohailiftikhar):

where tow lines will cut each other i will be the solution set of equeations...

OpenStudy (anonymous):

Oh okay so it should be (0,2) ?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

If you have a graph showing all of the equations, any solution to all of the equations will be a point at which all of the equations intersect. Is (2,0) such a point?

OpenStudy (whpalmer4):

For the same value of \(x\), both \(f(x)\) and \(g(x)\) must be equal for that value of \(x\) to be a solution.

OpenStudy (anonymous):

Oh so I just plug in 2 for x in both of my solutions and 0 for y?

OpenStudy (whpalmer4):

\[y = f(x) = |x-1|+1\]\[y = g(x) = 3x+2\]If we think \(x=2\) might be a solution, then \[f(2) = g(2)\]but \[f(2) = |2-1|+1 = 2\]and \[g(2) = 3(2) +2 = 8\] and those are not equal...

OpenStudy (whpalmer4):

Yes, you could also do it that way as a check: \[0 = f(2) = |2-1|+1\checkmark\]so far, so good, but let's try the other equation, which also must work: \[0 = g(2) = 3(2)+2 = 8\]Bzzt, wrong! So \((2,0\) is NOT a solution to that system of equations.

OpenStudy (whpalmer4):

sorry, \((2,0)\) is not a solution

OpenStudy (anonymous):

Okay thats what I thought. Thank you.